a)Find z such that 7.6% of the standard normal curve lies to the left of z. (Round your answer to two decimal places.)
z =

b)Find z such that 94.2% of the standard normal curve lies to the right of z. (Round your answer to two decimal places.)
z =
Answer #1

Solution:

Using standard normal table,

a ) P(Z < z) = 7.6%

P(Z < z) = 0.076

P(Z < - 1.43 ) = 0.076

z = - 1.43

b ) P(Z > z) = 94.2%

1 - P(Z < z) = 0.942

P(Z < z) = 1 - 0.942 = 0.058

P(Z < -1.57 ) = 0.058

z = -1.57